OpenStudy (s3a):

Integrating ∫√(4x - x^2) dx. Could someone please tell me how to start this one?

OpenStudy (turingtest):

complete the square is my first guess

OpenStudy (turingtest):

then trig sub

OpenStudy (anonymous):

follow @TuringTest you will get it

OpenStudy (kainui):

Trig sub is fun, we'll help you! Just give it your best try. @s3a

OpenStudy (anonymous):

well;\[\int\limits_{}^{}\sqrt{x}= \frac{ 2x ^{3/2} }{ 3 }\] so:\[\int\limits_{}^{}\sqrt{4x-x^2}=\frac{ (8x-2x^2)^{3/2} }{ 3 }\] and remember: \[x ^{3/2}=\sqrt{x^3}\] so:\[(8x-2x^2)^{3/2}=\sqrt{-8 x^6+96 x^5-384 x^4+512 x^3}\] going back to the original integral: \[\int\limits_{}^{}\sqrt{4x-x^2}=\frac{ \sqrt{-8 x^6+96 x^5-384 x^4+512 x^3} }{ 3 }\]

OpenStudy (anonymous):

i like using formulas for integration i hate u substitution and trig sub

OpenStudy (mathmale):

Seems to me that trig subst. is the way to go ... and that after completing the square. Let's focus on the quantity under the square root operator: 4x - x^2 = - (x^2 - 4x) Completing the square: - (x^2 - 4x + 4 - 4) Rewriting x^2 - 4x + 4, we get - ([x-2)^2 - 4), or 4 - (x-2)^2, or 2^2 - (x-2)^2. Then the integral becomes Int Sqrt( 2^2 - (x_2)^2 ) dx. Hint: if you have the situation Sqrt( a^2 - x^2), what is an appropriate substitution for x?

OpenStudy (s3a):

Actually, thank you everyone, but TuringTest's advice helped me get on the path I wanted. My problem now is that I am making a mistake somewhere in that path. Here is my work: http://www.tiikoni.com/tis/view/?id=ac56dfe Could someone please tell me where I went wrong?

OpenStudy (turingtest):

you should get it to be \(\sqrt{1+....}\) under the radical before substitution, so factor out that 4

OpenStudy (s3a):

Thanks, TuringTest (and everyone else). :) Here's my correct work: http://www.tiikoni.com/tis/view/?id=090cade